OpenStudy (anonymous):

what is the y intercept of the graph of 7x=3y-8

OpenStudy (anonymous):

for finding the y intercept you have to let x=0 and then solve for y

OpenStudy (anonymous):

put it in slope intercept form

OpenStudy (anonymous):

if i divid 3 from 0 ill get 0

OpenStudy (mindblast3r):

to put it in slope intercept form you gotta add and subtract. remember you gotta when you're doing it do it to both sides.

OpenStudy (mindblast3r):

and divide and multiply

OpenStudy (anonymous):

what? divide what

OpenStudy (anonymous):

subtract what?

OpenStudy (texaschic101):

sub in 0 for x 7x = 3y - 8 7(0) = 3y - 8 0 = 3y - 8 --- subtract 3y from both sides -3y = -8 -- now divide by -3 y = -8/-3 y = 8/3

OpenStudy (mindblast3r):